Please help me with this problem:
Find the slope and the y-intercept of the line
9x-4y=12
What is the slope?
What is the y-intercept?
Thank you so very much.

Answer by rfer(16322)

put in y intercept form
y=mx+b
-4y=-9x+12
y=9/4x-3
slope m=9/4
y int=-3